A baseball player has a lifetime batting average of 0.223. If, in a season, this player has 270 "at bats", what is the probability he gets 49 or more hits?

Probability of 49 or more hits = ?
Answer #1

Solution :

Given

p = 0.223

Number of trials = n = 270

We have to find P( X 49)

Here X follows binomial distribution with n = 270 and p = 0.223

P( X 49) = 1 - P(X 48) = 1 - 0.04081 = 0.9592 ( Using excel function =BINOMDIST(48,270,0.223,1) )

Please like if it helps me

Thank you so much